find the equation of a line that has a gradient of 3 and poassess through the point A(1,7)


1
Expert's answer
2021-03-07T10:12:27-0500

"y = kx + b" is an equation of a line, where "k" is the slope (or gradient) and "b" is the "y" -intercept;

"k = 3" by condition "\Rightarrow" "y = 3x + b" .

Since the line passes through the point "A (1,7)" , then we substitute the coordinates of the point into the equation: "7 = 3 \cdot 1 + b \Rightarrow b = 7-3 =4" "\Rightarrow" The equation of a line is "y = 3x + 4" .